OpenStudy (anonymous):

The seats at a local baseball stadium are arranged so that each row has five more seats than the row in front of it. If there are four seats in the first row, how many total seats are in the first 24 rows?

OpenStudy (anonymous):

Here is a good video on this topic: http://www.youtube.com/watch?v=Dj1JZIdIwwo

OpenStudy (anonymous):

\[(4*24) + \sum_{n = 1}^{24} 5n\]

OpenStudy (anonymous):

Partial sum = \[\frac{ n(a _{1} +a _{n})}{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 24(5(1)+5(24)) }{ 2 }\]

OpenStudy (anonymous):

\[12(5+120)\]

OpenStudy (anonymous):

So, \[\sum_{n = 1}^{24} 5n = 1500\]

OpenStudy (anonymous):

Now add (4*24) to 1500, and get your final answer: 1596